if you osve this,can you please break down the steps and explain what to do because i dont know what to do when you 3 equations like these:
A=11B
A+x=5(B+x)
A+x+5=3(B+x+5)

Answer by checkley77(12844)

First you replace A by 11B in the remaing 2 equations.
A+x=5(B+x)
11B+x=5(B+x)
11B+x=5B+5x
x-5x=5B-11B
-4x=-6B (EQUATION #1)
A+x+5=3(B+x+5)
11B+x+5=3(B+x+5)
11B+x+5=3B+3x+15
x-3x=-11B+3B+15
-2x=-8B+15 (EQUATION #2)
MULTIPLY THIS EQUATION BY -2 & ADD IT TO EQUATION #1.
4X=16B-30
-4x=-6B
------------------
0=10B-30
10B=30
B=30/10
B=3 ANSWER.
A=11*3
A=33 ANSWER.
-4x=-6*3
-4x=-18
x=-18/-4
X=4.5 ANSWER.
PROOF
33+4.5=5(3+4.5)
37.5=5*7.5
37.5=37.5